Determine the nature of the solutions of the equation 10t^2 - 8t = 0.
2) give exact and approximate solutions to thee decimal places x^2+2x+1=4
3) solve by copleting the square x^2+13/2x=13
4) give exact and approximate solutions to three decimal places (x-5)^2=18
Thank you,

Answer by ewatrrr(24785)

 
Hi,
10t^2 - 8t = 0 2t(5t-4)=0 t = 0,.8 (2-real solutions)
x^2+2x+1=4
x^2+2x-3=0 (x+3)(x-1)= 0 x = -3,1
x^2+13/2x=13
(x+13/4)^2 = 13 + 169/16 = 208/16 + 16*/169 = 377/16
x + 13/4 = ± sqrt(377)/4
x = -13/4 ± sqrt(377)/4 | x = -13/4 + sqrt(377/4),-13/4 - sqrt(377/4)
(x-5)^2=18
x-5 = ± sqrt(18)= ± sqrt(9*2)= ± 3sqrt(2)
x = 5 ± 3sqrt(2) | x = 5+3sqrt(2), 5-3sqrt(2) (exact solutions)
x = 5 ± 4.243 | x = .757, 9.243 (approximate solutions)
x^2 + 1/2x = 13
x(+1/4)^2 = 13 + 1/16 = 208/16 + 1/16 = 209/16
x = -1/4 +sqrt(209)/4, -1/4-sqrt(209)/4